OpenStudy (anonymous):

Factor 3x^2 - 15x - 42. 3(x + 2)(x - 7) 3(x - 2)(x + 7) 3(x - 7)(x + 2)

Parth (parthkohli):

\[\Large \color{purple}{\rightarrow 3(x^2 - 5x - 14) }\] Factor x^2 - 5x - 14 \[\Large \color{purple}{\rightarrow x^2 + 7x - 2x - 14 }\] \[\Large \color{purple}{\rightarrow x(x + 7) -2(x + 7) }\] \[\Large \color{purple}{\rightarrow (x - 2)(x + 7) }\] Answer would be ?

OpenStudy (anonymous):

second one

Parth (parthkohli):

Yes.

OpenStudy (callisto):

3x^2 - 15x - 42 = 3(x^2 - 5x -14) = 3(x-7)(x+2)

OpenStudy (anonymous):

so callisto u r saying the bottom one

Parth (parthkohli):

Oh yes, I'm wrong